# How Reliable is Language Model Micro-Benchmarking?

### v1

@@ 5.3 MDAD can interpret which model comparisons will be preserved
We have so far established that at small sizes, most micro-benchmarks can distinguish only those models whose performances differ greatly on the full benchmark. This holds when comparing target models across various sizes and training regimes. But what about when comparing a specific set of models that we expect might have more similar performances?
@@
Consider 32 instruction-tuned 8B-parameter models on MMLU-Pro. Most have accuracies between 27 and 40 on the full benchmark, yielding very low pairwise accuracy differences (Figure 5, left). MDAD computed from these 8B-parameter models (Figure 5, right) helps us understand when model comparisons are preserved on micro-benchmarks of various sizes. All micro-benchmarks have an MDAD of 5 or more when selecting 10 or 25 examples. These micro-benchmarks are not likely to reproduce the full benchmark’s ranking on the 51% of model comparisons that differ by at most 5 points of accuracy. When 1,000 examples are selected, most micro-benchmarks have an MDAD of 2. These can consistently rank more models, though they will still not be able to consistently rank the 21% of model pairs that differ by no more than 2 points of accuracy.
@@
Perlitz et al. (2024) observe that micro-benchmarks can often consistently predict the top-ranked target models, even when selecting few examples. MDAD offers a mechanism by which this occurs: top-ranked models often have high pairwise accuracy differences from many models, and micro-benchmarks often agree with the full benchmark when comparing very different models. That is, top-ranked models often differ from many models by more than a micro-benchmark’s MDAD. While all micro-benchmarks have high agreement with the full benchmark for the top-performing models once 25 or more examples are selected, they are less likely to agree with the full benchmark when comparing the models in the middle of the distribution that are closer in accuracy to each other (Figure 14, Appendix K). Figure 5 (right) shows that when selecting 25 examples, all methods achieve an MDAD of 5 or more. For the top-ranked model, two-thirds of all model comparisons are above this MDAD. For a model in the middle of the distribution, only one third of all pairwise comparisons are above this MDAD. Appendix K shows similar results when restricting comparisons to various other sets of models, like instruction-tuned 70B-parameter models.
Figure 6: MDAD is modestly higher on MMLU-Pro when predicting relative model performance on a held-out draw of the task (dashed lines) than when predicting relative performance on the full dataset used to select the micro-benchmarks (solid lines). See Appendix I for results on other datasets.
@@ 5.4 Micro-benchmarks generalize to new evaluation sets
One potential advantage of micro-benchmarks is that they may be able to exploit correlations between model predictions in order to predict how well models will do on fresh draws of the task. Or they could instead overfit their predictions to the specific examples they were selected from. We test whether either is the case by evaluating how well model comparisons on a micro-benchmark predict model comparisons on a held-out unseen set that was not used to construct the micro-benchmark. We find almost no difference in MDADs when predicting model performance on a held-out set using micro-benchmarks that select from entire benchmarks, indicating that these micro-benchmarking methods can effectively generalize to new draws of a dataset (full results in Appendix H). However, when selecting micro-benchmarks of subtasks individually, we do find that micro-benchmarks are slightly less able to predict model performance on fresh draws of the subtask, as evidenced by higher MDADs (Figure 6). Anchor Points experiences the least increase in MDAD when moving from the full benchmark to a fresh draw of the benchmark. Diversity and tinyBenchmarks experience larger increases. In all cases, MDAD is high. See Appendix I for full details and results.
6 Related Work
Efficient evaluation.
Benchmark datasets have been used to evaluate machine learning models for decades (Koch et al., 2021). Though benchmarks are not sufficient to fully map capabilities (Ethayarajh & Jurafsky, 2020; Kiela et al., 2021; Chiang et al., 2024), they remain an important tool for comparing models (Saxon et al., 2024). NLP benchmarks have grown ever-larger as LM capabilities have expanded (Srivastava et al., 2023; Liang et al., 2023). Micro-benchmarking methods select a small subset of a benchmark for evaluation with the goal of estimating a model’s performance on the full benchmark. This is done by exploiting correlations in a model’s predictions across examples (Fogliato et al., 2024b) and correlations across multiple models’ predictions (Vivek et al., 2024; Ye et al., 2023; Liu et al., 2023), by training instance difficulty models using Item Response Theory (Polo et al., 2024; Vania et al., 2021; Rodriguez et al., 2021), or even by deduplication (Gupta et al., 2025). We study the reliability of micro-benchmarks that summarize performance across an entire dataset, though other work focuses on estimating performance across subtasks (Fogliato et al., 2024a). Farther afield, active testing instead selects test instances in an online per-model manner (Kossen et al., 2021). Still other works use model-based proxies to select unlabeled examples for annotation (Tahan et al., 2025; Zouhar et al., 2025). Subsets of training data can be selected using gradients (Everaert & Potts, 2024; Xia et al., 2024; Engstrom et al., 2025) or information gain (Deb et al., 2025), though such methods have not yet been applied to evaluation data selection.
Evaluation reliability.
For classification, larger evaluation datasets yield more reliable model comparisons (Shalev-Shwartz & Ben-David, 2014; Dror et al., 2018; Card et al., 2020) and are more robust to some forms of dataset reuse (Yauney & Mimno, 2024). Card et al. (2020) use statistical power to estimate the minimum detectable effect size afforded by a benchmark, though this approach requires assumptions about generalization error and expected per-example agreement between model predictions. In contrast, our approach directly estimates the probability that pairwise model comparisons on a micro-benchmark agree with the full benchmark for a set of target models. Perlitz et al. (2024) also frame reliability as consistency over random choices in order to study the reliability of individual model error rate and model rankings over evaluation choices, but they do not identify which model comparisons are preserved by smaller datasets. Madaan et al. (2024) measure the variance in performance across random seeds and find that some micro-benchmarking methods increase variance. Our approach offers an interpretation of when variance impacts model comparisons. Other work considers the reliability afforded by LM generations for human evaluation (Ghosh et al., 2024; Boubdir et al., 2023).
7 Discussion
@@
We have investigated how well various micro-benchmarking methods reproduce model performance judgments from full benchmarks using the meta-evaluation measures of agreement and Minimum Detectable Ability Difference (MDAD). We find that when micro-benchmarks produce model rankings with high aggregate correlation with the rankings from full benchmarks, they cannot always consistently distinguish model pairs with similar performance. Once enough examples are selected to allow for distinguishing model pairs with similar performance, random sampling is competitive with other methods. Our meta-evaluation measures can guide micro-benchmarking method designers in building reliable and efficient model comparisons. While our experiments primarily focus on accuracy-based evaluation, there exist straightforward extensions of MDAD to other metrics, including ones used in open-ended generation.
@@
We release source code for reproducing all experiments, including all implementation details for datasets and micro-benchmarking methods. Appendix A describes the micro-benchmarking methods in detail, and Appendix C describes our experimental setup and infrastructure in greater detail. The GitHub repository also includes all intermediate micro-benchmarking results as well as the final results reported in figures throughout this paper.
@@ Acknowledgments
Appendix A Details of Micro-Benchmarking Methods
@@
Appendix B Different Agreement Thresholds for MDAD
Figure 7 shows a comparison of 0.7, 0.8, and 0.9 as the threshold of agreement in Equation 5. All of our other experiments use a threshold of 0.8.
Figure 7: MDADs when using different thresholds for agreement are qualitatively similar. The middle row of MDAD panels are the same as in Figure 4. Error bars represent 95% bootstrap confidence intervals over 50 trials.
Appendix C Full experimental setup and computing infrastructure
For most experiments, we construct micro-benchmarks using 300 source models and evaluate them using 50 target models. We average all evaluation metrics over 50 runs of random partitions into source and target models. For the experiments in Section 5.2 and Appendix K, we choose 300 random source models and evaluate with a fixed set of (non-overlapping) target models. Experiments were implemented using NumPy (Harris et al., 2020). We report calculated MDADs to the nearest 0.5.
@@ Appendix C Full experimental setup and computing infrastructure
Appendix D Quantitative Error Analysis of MDAD
@@
Appendix E Different Bucket Resolutions for MDAD
@@
Appendix F Full results
@@
Appendix G MDAD Can Analyze Methods that Select Entire Subtasks
Throughout this paper, we evaluate micro-benchmarking methods that select examples without regard to which subtasks of the original benchmark they are from. But our methods and proposed meta-evaluation measures generalize to other kinds of micro-benchmarking methods as well. For example, BenTo is a micro-benchmarking method that selects whole subtasks for smaller evaluation sets by estimating task transferability (Zhao et al., 2024). When selecting a micro-benchmark from MMLU, BenTo selects 801 examples in three subtasks of MMLU for micro-benchmarking. We calculate MDAD and other meta-evaluation measures for this selected subset of examples, as well as uniform random sampling for selecting 801 examples (Table 3). Both BenTo and Random achieve very similar results.
@@ Appendix G MDAD Can Analyze Methods that Select Entire Subtasks
Appendix H Generalizing to New Task Draws when Selecting from Entire Benchmarks
@@
Appendix I Generalizing to New Task Draws when Selecting from Subtasks Separately
@@
Appendix J Increasing Number of Source Models has Modest Effect
@@
Appendix K Comparing specific sizes of models

### v2

@@ 5.3 MDAD can interpret which model comparisons will be preserved
We have so far established that at small sizes, most micro-benchmarks can distinguish only those models whose performance differ greatly on the full benchmark. This holds when comparing target models across various sizes and training regimes. But what about when comparing a specific set of models that we expect might have more similar performances? We consider a case study of 32 instruction-tuned 8B-parameter models on MMLU-Pro.
Most models have accuracies between 27 and 40 on the full benchmark, yielding very low pairwise accuracy differences (Figure 5, left). MDAD computed from these 8B-parameter models (Figure 5, right) helps us understand when model comparisons are preserved on micro-benchmarks of various sizes. All micro-benchmarks have an MDAD of 5 or more when selecting 10 or 25 examples. These micro-benchmarks are not likely to reproduce the full benchmark’s ranking on the 51% of model comparisons that differ by at most 5 points of accuracy. When 1,000 examples are selected, most micro-benchmarks have an MDAD of 2. These can consistently rank more models, though they will still not be able to consistently rank the 21% of model pairs that differ by no more than 2 points of accuracy.
@@
Perlitz et al. (2024) observe that micro-benchmarks can often consistently predict the top-ranked target models, even when selecting few examples. MDAD offers a mechanism by which this occurs: top-ranked models often have high pairwise accuracy differences from many models, and micro-benchmarks often agree with the full benchmark when comparing very different models. That is, top-ranked models often differ from many models by more than a micro-benchmark’s MDAD. While all micro-benchmarks have high agreement with the full benchmark for the top-performing models once 25 or more examples are selected, they are less likely to agree with the full benchmark when comparing the models in the middle of the distribution that are closer in accuracy to each other (Figure 14, Appendix L). Figure 5 (right) shows that when selecting 25 examples, all methods achieve an MDAD of 5 or more. For the top-ranked model, two-thirds of all model comparisons are above this MDAD. For a model in the middle of the distribution, only one third of all pairwise comparisons are above this MDAD. Appendix L shows similar results when restricting comparisons to various other sets of models, like instruction-tuned 70B-parameter models.
@@ 5.4 Micro-benchmarks generalize to new evaluation sets
One potential advantage of micro-benchmarks is that they may be able to exploit correlations between model predictions in order to predict how well models will do on fresh draws of the task. Or they could instead overfit their predictions to the specific examples they were selected from. We test whether either is the case by evaluating how well model comparisons on a micro-benchmark predict model comparisons on a held-out unseen set that was not used to construct the micro-benchmark. We find almost no difference in MDADs when predicting model performance on a held-out set using micro-benchmarks that select from entire benchmarks, indicating that these micro-benchmarking methods can effectively generalize to new draws of a dataset (full results in Appendix I). However, when selecting micro-benchmarks of subtasks individually, we do find that micro-benchmarks are slightly less able to predict model performance on fresh draws of the subtask, as evidenced by higher MDADs (Figure 6). Anchor Points experiences the least increase in MDAD when moving from the full benchmark to a fresh draw of the benchmark. Diversity and tinyBenchmarks experience larger increases. In all cases, MDAD is high. See Appendix J for full details and results.
Figure 6: MDAD is modestly higher on MMLU-Pro when predicting relative model performance on a held-out draw of the task (dashed lines) than when predicting relative performance on the full dataset used to select the micro-benchmarks (solid lines). See Appendix J for results on other datasets.
6 Discussion and Conclusion
@@
We have investigated how well various micro-benchmarking methods reproduce model performance judgments from full benchmarks using the meta-evaluation measures of agreement and Minimum Detectable Ability Difference (MDAD). We find that when micro-benchmarks produce model rankings with high aggregate correlation with the rankings from full benchmarks, they cannot always consistently distinguish model pairs with similar performance. Once enough examples are selected to allow for distinguishing model pairs with similar performance, random sampling is competitive with other methods. Our meta-evaluation measures can guide micro-benchmarking method designers in building reliable and efficient model comparisons. While our experiments primarily focus on accuracy-based evaluation, there exist straightforward extensions of MDAD to other metrics, including ones used in open-ended generation. In future work, we plan to explore using MDAD to directly guide data selection strategies.
@@
We release source code for reproducing all experiments, including all implementation details for datasets and micro-benchmarking methods, on GitHub. Appendix B describes the micro-benchmarking methods in detail, and Appendix D describes our experimental setup and infrastructure in greater detail. The GitHub repository also includes all intermediate micro-benchmarking results as well as the final results reported in figures throughout this paper.
@@ Acknowledgments
Appendix A Additional Related Work
Efficient evaluation.
Benchmark datasets have been used to evaluate machine learning models for decades (Koch et al., 2021). Though benchmarks are not sufficient to fully map capabilities (Ethayarajh & Jurafsky, 2020; Kiela et al., 2021; Chiang et al., 2024), they remain an important tool for comparing models (Saxon et al., 2024). NLP benchmarks have grown ever-larger as LM capabilities have expanded (Srivastava et al., 2023; Liang et al., 2023). Micro-benchmarking methods select a small subset of a benchmark for evaluation with the goal of estimating a model’s performance on the full benchmark. This is done by exploiting correlations in a model’s predictions across examples (Fogliato et al., 2024b) and correlations across multiple models’ predictions (Vivek et al., 2024; Ye et al., 2023; Liu et al., 2023), by training instance difficulty models using Item Response Theory (Polo et al., 2024; Vania et al., 2021; Rodriguez et al., 2021), or even by deduplication (Gupta et al., 2025). We study the reliability of micro-benchmarks that summarize performance across an entire dataset, though other work focuses on estimating performance across subtasks (Fogliato et al., 2024a). In contemporaneous work, Zhang et al. (2025) give complementary results on the efficacy of random sampling for predicting individual model accuracy. They find that a regression-based extension of random sampling has competitive mean estimation error with other micro-benchmarking methods when selecting 50 data points. They also study extrapolating from lower-accuracy source models to high-accuracy target models, finding that random sampling and a strategy based on augmented inverse propensity weighting achieve lower mean estimation error than other methods.
Farther afield, active testing instead selects test instances in an online per-model manner (Kossen et al., 2021). Still other works use model-based proxies to select unlabeled examples for annotation (Tahan et al., 2024; Zouhar et al., 2025). Subsets of training data can be selected using gradients (Everaert & Potts, 2024; Xia et al., 2024; Engstrom et al., 2025) or information gain (Deb et al., 2025), though such methods have not yet been applied to evaluation data selection.
Evaluation reliability.
For classification, larger evaluation datasets yield more reliable model comparisons (Shalev-Shwartz & Ben-David, 2014; Dror et al., 2018; Card et al., 2020) and are more robust to some forms of dataset reuse (Yauney & Mimno, 2024). Card et al. (2020) use statistical power to estimate the minimum detectable effect size afforded by a benchmark, though this approach requires assumptions about independence of examples, generalization error, and expected per-example agreement between model predictions. In contrast, our approach directly estimates the probability that pairwise model comparisons on a micro-benchmark agree with the full benchmark for a set of target models. Perlitz et al. (2024) also frame reliability as consistency over random choices in order to study the reliability of individual model error rate and model rankings over evaluation choices, but they do not identify which model comparisons are preserved by smaller datasets. Madaan et al. (2024) measure the variance in performance across random seeds and find that some micro-benchmarking methods increase variance. Our approach offers an interpretation of when variance impacts model comparisons. Other work considers the reliability afforded by LM generations for human evaluation (Ghosh et al., 2024; Boubdir et al., 2023).
Appendix B Details of Micro-Benchmarking Methods
@@
Appendix C Different Agreement Thresholds for MDAD
Figure 7 shows a comparison of 0.7, 0.8, 0.9, and 0.95 as the threshold of agreement in Equation 5. All of our other experiments use a threshold of 0.8. Agreement between a micro-benchmark and the full benchmark is higher for larger differences in model performance As the agreement threshold for MDAD increases, the main effect is that MDAD also increases. All methods tend to experience this MDAD increase, so the overall results for different thresholds are qualitatively similar.
Figure 7: MDADs when using different thresholds for agreement are qualitatively similar. The second row of MDAD panels with a 0.8 threshold are the same as in Figure 4. Error bars represent 95% bootstrap confidence intervals over 50 trials.
Appendix D Full experimental setup and computing infrastructure
For most experiments, we construct micro-benchmarks using 300 source models and evaluate them using 50 target models. We average all evaluation metrics over 50 runs of random partitions into source and target models. For the experiments in Section 5.2 and Appendix L, we choose 300 random source models and evaluate with a fixed set of (non-overlapping) target models. Experiments were implemented using NumPy (Harris et al., 2020). We report calculated MDADs to the nearest 0.5.
@@ Appendix D Full experimental setup and computing infrastructure
Appendix E Quantitative Error Analysis of MDAD
@@
Appendix F Different Bucket Resolutions for MDAD
@@
Appendix G Full results
@@
Appendix H MDAD Can Analyze Methods that Select Entire Subtasks
Throughout this paper, we evaluate micro-benchmarking methods that select examples without regard to which subtasks of the original benchmark they are from. But our methods and proposed meta-evaluation measures generalize to other kinds of micro-benchmarking methods as well. For example, BenTo is a micro-benchmarking method that selects whole subtasks for smaller evaluation sets by estimating task transferability (Zhao et al., 2025). When selecting a micro-benchmark from MMLU, BenTo selects 801 examples in three subtasks of MMLU for micro-benchmarking. We calculate MDAD and other meta-evaluation measures for this selected subset of examples, as well as uniform random sampling for selecting 801 examples (Table 4). Both BenTo and Random achieve very similar results.
@@ Appendix H MDAD Can Analyze Methods that Select Entire Subtasks
Appendix I Generalizing to New Task Draws when Selecting from Entire Benchmarks
@@
Appendix J Generalizing to New Task Draws when Selecting from Subtasks Separately
@@
Appendix K Increasing Number of Source Models has Modest Effect
@@
Appendix L Comparing specific sizes of models